A bus leaves a station at 1P.M. traveling west at an average rate of 44 mi/h. One hour later a second bus leaves the same station, traveling east at a rate of 48 mi/h. At what time will the two buses be 274 miles apart?
Answer by checkley75(3666)

44x+48(x-1)=274
44x+48x-48=274
92x=274+48
92x=322
x=322/92
x=3.5 hours later the buses were 274 miles apart.